# Incremental Joint Learning of Depth, Pose and Implicit Scene Representation on Monocular Camera in Large-scale Scenes

### v1

@@ III-D Incremental Joint Learning
IV Experiments
Baselines Methods and Public Datasets. The main baseline methods we compared are BARF [16], NoPe-NeRF [22], and LocalRF [23]. All of them are joint learning methods of camera poses and implicit scene representation. We evaluate our method on the Tanks and Temples dataset [21] and Static Hikes dataset [23]. The Tanks and Temples dataset features slow-moving image sequences spanning a few meters in both indoor and outdoor scenes. We also utilize the Static Hikes indoor dataset, which includes hand-held sequences with longer camera trajectories, testing scalability and pose estimation robustness over ranges typically exceeding 30 meters. We strictly follow the experimental settings of our baseline to partition the training set and the test set.
Implementation Details. We present our implementation details. We select Adam optimizer[52] ($\beta=(0.9,0.999)$) for scene representation and camera tracking optimization. The color loss weighting is $\lambda_{1}=0.25$, $\lambda_{2}=0.1$ for depth and $\lambda_{3}=1.0$ for pose estimation, $\lambda_{4}=1.0$, $\lambda_{5}=1.0$ for optical flow. All training and evaluation experiments are conducted on a single NVIDIA RTX 3090 GPU.
@@ IV-A Experimental Results
Scene Reconstruction Results. We use Peak Signal-to-noise Ratio (PSNR), Structural Similarity Index Measurement (SSIM), and Learned Perceptual Image Patch Similarity (LPIPS) to evaluate view synthesis results. In Table I, we present the scene reconstruction results on Tanks and Temples datasets. Our method outperforms all the baselines by a large margin. The scene reconstruction results on other datasets are shown in Table III. Our method outperforms other methods in scene representation. The incremental implicit scene representation method significantly enhances the scene representation ability and performs well on large-scale and long-sequence datasets. The qualitative results are shown in Fig. 3,4,5. In Fig. 3, We present the view synthesis results on different sequences Ignatius, Horses, Church in Tanks and Temples dataset [21]. Our method achieves higher-quality image reconstruction results. In Fig. 4, these images correspond to different timesteps (90,390,900) from up to down. We place the PSNR values of each scene in the bottom right corner. In contrast, our incremental scene representation method achieves high-quality view synthesis results compared with NoPe-NeRF and LocalRF. In Fig. 5, we show the image reconstruction results on our own dataset. Compared with other joint learning methods, our method achieves the most high-fidelity novel view results in both outdoor and indoor scenes.
Pose Estimation Results. The pose estimation results are shown in Table. II and Table. III. We use Absolute Trajectory Error (ATE), and Relative Pose Error (RPE) as the pose estimation metrics. Our pose estimation network performs better than other existing methods. The FBA method performs re-projection optimization in the high-dimensional feature level, which avoids various disturbances and noise present at the image level. The FBA network enhances the accuracy and robustness of the pose estimation in large-scale scenes. We also present the depth estimation results in Fig. 6. Compared to our method, a tree is missing from the LocalRF estimated depth, as shown in red blocks, and the depth at both bushes and trees in NoPe-NeRF is inaccurate, as shown in black blocks. Our estimated depth images are more accurate and sharper, as shown in green blocks.
@@
TABLE IV: Ablation studies of different modules on Tanks and Temples datasets [21].
Tanks and Temples [21]
[…]
(a) w/o FBA-net 27.15 0.87 0.10 0.017 0.157 0.143
(b) w/o PO 25.87 0.84 0.15 0.016 0.152 0.141
(c) w/o $L_{d}$ 26.75 0.86 0.11 0.017 0.143 0.137
Ours 29.11 0.90 0.09 0.013 0.139 0.133
@@ IV-B Ablation Study
In this section, we conduct sufficient ablation studies to verify the effectiveness of our designed network. We show the ablation results in Table IV. (a) is the network without a FBA network. (b) is the network without progressive optimization. (c) is the network without depth loss. For the FBA network, we remove it in Table IV. (a), and replace it with traditional pose optimization. We can see that this network improves the accuracy of the pose estimation results. For the progressive optimization network, we replace it with global optimization. We can see that the progressive network significantly improves the scene representation and pose estimation. We also present the render results between using a single global radiance field and using multiple radiance fields in Fig. 7. This indicates the importance of our incremental scene representation method. In Table (c), we remove the depth loss and the depth network is not jointly optimized with other networks. The results indicate that the depth loss and joint optimization strategy are effective for scene representation and pose estimation.

### v2

@@ IV-A Experimental Results
Scene Reconstruction Results. We use Peak Signal-to-noise Ratio (PSNR), Structural Similarity Index Measurement (SSIM), and Learned Perceptual Image Patch Similarity (LPIPS) to evaluate view synthesis results. In Table I, we present the scene reconstruction results on Tanks and Temples datasets. Our method outperforms all the baselines by a large margin. In Tab. III, we compare our method with other view synthesis methods with COLMAP pose, which shows the proves the advantages of our joint learning system in large-scale scenes. The scene reconstruction results on other datasets are shown in Table IV. Our method outperforms other methods in scene representation. The incremental implicit scene representation method significantly enhances the scene representation ability and performs well on large-scale and long-sequence datasets. The qualitative results are shown in Fig. 3,4,5. In Fig. 3, We present the view synthesis results on different sequences Ignatius, Horses, Church in Tanks and Temples dataset [8]. Our method achieves higher-quality image reconstruction results. In Fig. 4, these images correspond to different timesteps (90, 390, 900) from up to down. We place the PSNR values of each scene in the bottom right corner. In contrast, our incremental scene representation method achieves high-quality view synthesis results compared with NoPe-NeRF and LocalRF. In Fig. 5, we show the image reconstruction results on our own dataset. Compared with other joint learning methods, our method achieves the most high-fidelity novel view results in both outdoor and indoor scenes.
Pose Estimation Results. The pose estimation results are shown in Table. II and Table. IV. We use Absolute Trajectory Error (ATE), and Relative Pose Error (RPE) as the pose estimation metrics. Our pose estimation network performs better than other existing methods. The FBA method performs re-projection optimization in the high-dimensional feature level, which avoids various disturbances and noise present at the image level. The FBA network enhances the accuracy and robustness of the pose estimation in large-scale scenes. We also present the depth estimation results in Fig. 6. Compared to our method, a tree is missing from the LocalRF estimated depth, as shown in red blocks, and the depth at both bushes and trees in NoPe-NeRF is inaccurate, as shown in black blocks. Our estimated depth images are more accurate and sharper, as shown in green blocks.
@@
TABLE V: Ablation studies of different modules on Tanks and Temples datasets [8].
Tanks and Temples
PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ ATE $\downarrow$ RPEr $\downarrow$ RPEt $\downarrow$
(a) w/o Dep. 26.75 0.86 0.11 0.016 0.063 0.098
(b) w/o ISR 26.57 0.84 0.15 0.015 0.065 0.101
(c) w/o FBA 27.15 0.87 0.10 0.016 0.068 0.103
(d) w COLMAP 28.17 0.88 0.10 0.014 0.051 0.081
Ours 29.04 0.90 0.09 0.012 0.046 0.077
@@ IV-B Ablation Study
In this section, we conduct sufficient ablation studies to verify the effectiveness of our designed network. We show the ablation results in Table V. (a) we remove the depth estimation and depth prior. (b) is the network without incremental scene representation. (c) is the network without feature-metric bundle adjustment. (d) is our method using pose from COLMAP. In Table (a), we remove the depth loss and the depth network is not jointly optimized with other networks. The results indicate that the depth loss and joint optimization strategy are effective for scene representation and pose estimation. (b) For the incremental scene representation, we replace it with global optimization. We can see that the incremental scene representation significantly improves the scene representation and pose estimation. We also present the render results between using a single global radiance field and using multiple radiance fields in Fig. 7. This indicates the importance of our incremental scene representation method. (c) We remove the feature-metric bundle adjustment and replace it with traditional pose optimization. We can see that this network improves the accuracy of the pose estimation results. (d) We use the camera pose from COLMAP and fix the camera pose throughout the entire optimization. This demonstrates that our joint optimization method outperforms the original method based on COLMAP.